• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name jpchen

    Log entry time 18:35:37 on August14,2001

    Entry number 70727

    keyword=Laser wave length checked and adjusted

    Check and sdjusted laser wavelength: (With help from Dave, Kevin and Mike)
    Before starting checking, three longitudinal lasers were on.
    All longitudinal lasers were pretty good. Lasers #6, #7, do not need adjustment.
    #5 adjusted slightly, temperature was set/read=18.0/17.9 adjusted to 18.1/17.9.

    Transverse lasers all needed adjustment: (they were off before checking)

    Laser T( Set/Read) After Adjustment
    #1 13/13 14/14
    #2 20.7/21.0 22.1/22.1
    #3 26.2/26.2 28.0/27.2
    #4 15.0/14.8 15.4/15.0

    Then changed to transverse pumping with lasers #1,2,4 on for about 1.5 hours.
    Check/and adjusted again:
    #1 and #4 unchanged
    #2, from 22.1/22.1 to 22.4/22.3
    #3, from 28.0/27.5 to 27.9/27.4

    Hope these will also help with the transverse pumping efficiency.